• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name hyde

    Log entry time 10:59:50 on October 06, 2010

    Entry number 333791

    keyword=Compton eDet noise

    One channel on right and left side of Compton detector is monitored by analog card and scope in counting house middle room.

    Turned on eDet electronics and Si HV to 100 (DAC? HV?) (most recent runs with eDet on, Compton logbook says 255)

    Noise value terminated into 50 Ohm at end of long cable is approx 1.6mV rms.

    Will try tomorrow with histogramming scope